Assume that annual interest rates are 6 percent in the United States and 5 percent in Turkey. An FI can borrow (by issuing CDs) or lend (by purchasing CDs) at these rates. The spot rate is $0.6553/Turkish lira (TL).

a. If the forward rate is $0.6735/TL, how could the bank arbitrage using a sum of $6 million? What is the spread earned? (Do not round intermediate calculations. Round your answer to 4 decimal places. (e.g., 32.1616))
b. At what forward rate is this arbitrage eliminated? (Do not round intermediate calculations. Round your answer to 5 decimal places. (e.g., 32.16161))
Answer #1

a. Bank can earn by borrowing the same amount in USA and Investing the same at TL.
It will earn an arbitrage profit of = $2,36,639.707

b. The forward rate should be $0.6491/TL

As per cip; Es itt HOTL rs = 0.6735 x (1-06) - 0.6553 ro = 9.235% (should be) re- 5.1. (Actual) Borrow in $& Invest in TL ③As per Covered Interest rate Panity. E = 1+84 Es Forward rate (should be) so spot rate ($0.6553/TL) TB = interest rate in USA